Assume that a ball is thrown vertically upward with initial velocity fo 96 ft per second. The distance s(t) (in feet) of the ball from the ground after t seconds is s(t)=96t-16t^2.
a. at what instant will it be back at the ground level?
b. for what time interval is the ball more than 112ft above the ground?

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE WHICH IS SIMILAR TO YOURS AND TRY.
----------------------------------------------------------
1. Assume that a ball is thrown vertically upward with initial velocity of
> 96 ft. per second. The distance s(t) (in feet) of the ball from the ground
> after t seconds is s(t) = 96t-16t^2.
>
> a. at what instant will it be back at the ground level?
AT GROUND S=0=96T-16T^2=16T(T-9)
HENCE T=0 ..THAT IS BEFORE IT IS THROWN UP AND
T=9...SO AFTER 9 SECS.IT WILL HIT THE GROUND ON FALL
> b. for what time interval is the ball more than 128 ft. above the ground.
> S=128=96T-16T^2.DIVIDING BY 16
T^2-6T+8=0=T^2-4T-2T+8=T(T-4)-2(T-4)=(T-4)(T-2)=0
T=2 SECS ON THE UPWARD TRAVEL AND
T=4 SECS ON THE DOWNWARD TRAVEL.HENCE IT WILL BE OVER 128 FT.HIGH
DURING T=2 TO 4 SECS.